Use two equations in two variables to solve the application.
Students can buy tickets to a basketball game for $2. The admission for nonstudents is $3. If 300 tickets are sold and the total receipts are $670, how many student tickets are sold?

Answer by ikleyn(52792)

.
Let x be the number of students and y be the number of others.


 x +  y = 300    (1)
2x + 3y = 670    (2)


To solve the system, multiply eq(1) by 2 (both sides). Keep eq(2) as is.


2x + 2y = 600    (1')
2x + 3y = 670    (2')


Subtract eq(1') from eq(2')  (both sides).  You will get


     3y - 2y = 670-600

        y    =  70.


Then from eq(1),  x = 300-70 = 230.


Answer.  230 students tickets were sold.

Solved.